Prove cos 5-sin 25 =sin 35 immediately correct answer

Grade:12th pass

1 Answers

Devang
5 years ago
                  L.H.S = cos5 – sin 25
                            = sin 95 – sin 25                                                             [ sin(90+5) = cos5 ]
                            = 2 * cos (95+25)/2 * sin (95-25)/2
                            = 2 * cos60 * sin35
                            =2 * ½ * sin35                                                                 (cos60 = ½ )
                            = sin35
                           =  R.H.S
hence      L.H.S  = R.H.S